can someone help me with this problem the perimeter of a triangle is 51 meters the second is twice as long as the first side the third side is 3 meters less than three times the first what are the lengths of the sides of the triangle give three answers seperated by commas
Answer by rfer(16322)

x+2x+3x-3=51
6x=51+3
x=54/6
x=9
2x=18
3x-3=24
9m, 18m, 24m